# Localization behavior in a Hermitian and non-Hermitian Raman lattice

### v1

@@ Abstract
We introduce a flexible Raman lattice system for alakaline-earth like atoms to investigate localization behaviors in a quasi-periodic lattice with controllable non-Hermiticity. We demonstrate that critical phases and mobility edges can arise by adjusting spin-dependence of the incommensurate potentials in the Hermitian regime. With non-Hermiticity introduced by spin-selective atom loss, we find the localization behaviour in this system can be suppressed by dissipation. Our work provides insights into interplay between quasi-periodicity and non-Hermitian physics, offering a new perspective on localization phenomena.
I I. Introduction
@@
In the presence of a disorder potential, quantum systems can exhibit exponential localization of wave functions induced by wave interference, a phenomenon known as Anderson localization Anderson (1958). Ultracold atom systems, with their exceptional purity and high controllability, enable detailed exploration of localization phenomena. Recent experiments have investigated Anderson localization in optical lattices with either random disorder Billy et al. (2008) or incommensurate quasi-periodic structures Roati et al. (2008); Lüschen et al. (2018); Xiao et al. (2021).
@@
For ultracold atoms, quasi-periodic optical lattices are created by superimposing two optical lattices with incommensurate wavelengths. In such setups, a notable transition from extended to localized states, commonly referred to as the Anderson transition, can manifest even within one-dimensional configurations Grempel et al. (1982); Aulbach et al. (2004), contrary to uncorrelated disordered systems where such transitions are expected to occur only in dimensions higher than two according to the scaling theory Thouless (1974); Abrahams et al. (1979). These systems also allow for the engineering of mobility edges-boundaries that separate coexisting localized and extended states at discrete energy levels Sarma et al. (1988); Evers and Mirlin (2008); Biddle and Das Sarma (2010); Lüschen et al. (2018); An et al. (2018, 2021). Recent theoretical studies suggest even more complex behavior, including intermediate critical phases and hybrid phases where localized, extended, and critical states coexist in specific one-dimensional quasi-periodic models Wang et al. (2020, 2022).
@@
Recent years have seen growing interest in non-Hermitian physics El-Ganainy et al. (2018); Ashida et al. (2021) across multiple research domains, including electrical circuits Helbig et al. (2020); Ezawa (2019), optics and photonics Lee et al. (2009); Xiao et al. (2020); Zhou et al. (2018); Weidemann et al. (2020), optomechanics Patil et al. (2022), NV centers Zhang et al. (2021); Yu et al. (2022), trapped ions Cao et al. (2023); Zhang et al. (2025), superconducting qubits Chen et al. (2021) and ultracold atom systems Li et al. (2019); Liang et al. (2022); Ren et al. (2022). Non-Hermitian systems display unique features such as exceptional points or rings Zhen et al. (2015); Ren et al. (2022); Liu et al. (2025), novel topological invariants, and the non-Hermitian skin effect Liang et al. (2022); Zhao et al. (2025). The Hatano-Nelson model, a fundamental non-Hermitian topological model, combines asymmetric hopping amplitudes with an on-site random potential Hatano and Nelson (1996). However, a comprehensive understanding of the interplay between disorder and non-Hermiticity, particularly the influences of dissipation on localization behavior, remains largely unexplored Zeng et al. (2017); Zhang et al. (2020); Harter et al. (2016). Furthermore, experimental methods for studying this relationship between disorder and non-Hermiticity remain limited.
@@
In this paper, we introduce a flexible Raman lattice scheme Liu et al. (2013) tailored for 173Yb atoms Song et al. (2018, 2019) or alkaline-earth metal atoms Liang et al. (2023) to investigate localization behaviors. The Raman lattice configuration induces quasi-periodicity using two optical lattices with incommensurate wavelengths, with the weak lattice potential akin to an incommensurate Zeeman potential. By adjusting the frequency of the incommensurate lattice close to the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, the spin-dependent nature of the lattice potential may lead to the emergence of critical phases or mobility edges within the system. Finally, we generalize the Raman lattice with incommensurate potentials into the non-Hermitian regime with dissipation induced by atom loss Zhao et al. (2025). The analysis reveals that dissipation exerts a significant influence on the localization behaviors within the system, offering a platform to investigate the intricate interplay between disorder potential and non-Hermitian physics.
@@
The subsequent sections of this paper are structured as follows: Sec.II presents the model Hamiltonian of the flexible Raman lattice and outlines its experimental realization scheme. In Sec.III-V, we delve into the impact of the spin-dependency of the incommensurate lattice and dissipation within the system. Our focus will encompass three distinct scenarios: a completely spin-dependent incommensurate lattice, a partially spin-dependent incommensurate lattice, and an incommensurate lattice coupled with dissipation. Finally, Sec.VI offers a concise summary of the findings and conclusions derived from this investigation.
II II. Model Hamiltonian
@@
$H=H_{0}+\sum_{i,\sigma=\uparrow,\downarrow}(\delta_{\sigma}^{i}+i\gamma_{\sigma})n_{i,\sigma}$ (1)
@@
where $\delta^{i}_{\sigma}=M_{z,\sigma}\cos{\left(2\pi\beta i\right)}$, with $M_{z,\sigma}\propto V_{s,\sigma}$ and $\gamma_{\sigma}$ denotes the incommensurate Zeeman potential strength and dissipation strength for spin $\sigma=\uparrow,\downarrow$, respectively. $n_{i,\sigma}=c^{\dagger}_{i,\sigma}c_{i,\sigma}$ is the particle number operator, $\beta$ is an irrational number and:
$H_{0}$ $=-t_{0}\sum_{i,\pm,\sigma}c_{i,\sigma}^{\dagger}c_{i\pm 1,\sigma}$ (2)
[…]
@@
Here, $t_{0}$ ($t_{so}$) represents the spin-conserved (spin-flip) hopping strength between neighboring sites. The Hamiltonian renders a 1D AIII class topological insulator if replacing the incommensurate Zeeman potential with a constant uniform Zeeman potential Song et al. (2018); Liu et al. (2013); Wang et al. (2018); Zhang and Liu (2018).
@@
Figure 1: Experimental realization of critical phase in optical Raman lattice with 173Yb atoms. (a) Experimental setup of the optical Raman lattice consists of a standard 1D optical Raman lattice $E_{1}$, $E_{3}$ with another spin-dependent incommensurate lattice $E_{21}$ and $E_{22}$. (b) Schematic energy diagram with relevant transitions. (c) The incommensurate lattice induces a spin-dependent offset $V_{2,\sigma}(x)$ to the standard optical Raman lattice potential. (d) For the implementation of a perfect spin-dependent incommensurate lattice, the laser detuning should be set approximately 1.9 GHz blue-detuned from the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition.
@@
This Hamiltonian can be experimentally implemented through the configuration outlined in Fig. 1(a). To establish the conventional optical Raman lattice potential Liu et al. (2013); Wu et al. (2016); Song et al. (2018), a standing-wave beam $E_{1}$ with $y$ polarization and a free-running beam $E_{3}$ with z polarization are employed. The single-photon detuning of these two beams can be finely tuned to approximately 11 GHz blue-detuned from the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, as depicted in Fig. 1(b). In this setup, the primary lattice, characterized by a wave vector of $k_{0}=2\pi/\lambda$, exhibits nearly spin-independent lattice potential $V_{1,\sigma}(x)=V_{p,\sigma}\cos^{2}{k_{0}x}$ with $V_{p,\uparrow}/V_{p,\downarrow}\sim 1.01$, where $V_{p,\uparrow}/V_{p,\downarrow}$ denotes the primary lattice depth for spin up or spin down atoms. In the subsequent discussion, we will neglect the spin-dependent effects of the primary lattice potential ($V_{p,\uparrow}=V_{p,\downarrow}=V_{p}$). Moreover, two additional beams propagating at an angle of 52 degrees relative to the x direction are utilized to induce a spin-dependent incommensurate lattice featuring a wave vector of $k_{1}=k_{0}\cos{\theta}$, giving rise to a secondary incommensurate Zeeman potential $V_{2,\sigma}=V_{s,\sigma}\cos^{2}{k_{1}x}$ with $V_{s,\sigma}$ representing the incommensurate lattice depth, as illustrated in Fig. 1(c). The spin-dependence of this incommensurate lattice can be controlled through the detuning $\Delta$ of this lattice beam with respect to the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition. When $\Delta\sim 1.9~{}\text{GHz}$, the incommensurate lattice becomes entirely spin-dependent with $M_{z,\uparrow}/M_{z,\downarrow}=-1$ (Fig. 1(d)). To achieve a complete spin-independent configuration, the detuning of the incommensurate lattice can be set larger than 13GHz. The Raman coupling in this setup can be denoted by $\mathcal{M}(x)=M_{R}\cos{k_{0}x}$. The introduction of dissipation into this Raman lattice setup can be realized through adding a nearly resonant loss beam with wavelength close to ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition Zhao et al. (2025).
@@
[…]
III III. Localization behavior in the Hermitian regime
Except for the localized phase, recent theoretical investigations have predicted the existence of a third fundamental phase, called critical phases, positioned between the localized and extended phases within such a spin-orbit coupled optical lattice featuring an incommensurate Zeeman potential Wang et al. (2020); Song et al. (2018); Zhao et al. (2025). Noteworthy features of these critical phases encompass critical spectral statistics Geisel et al. (1991); Machida and Fujita (1986); Bertrand and García-García (2016), multifractal properties of wave functions Halsey et al. (1986); Mirlin et al. (2006); Dubertrand et al. (2014), and dynamical evolutions Hisashi and Shuji (1988); Ketzmerick et al. (1997); Larcher et al. (2009). Moreover, recent theoretical investigations have also unveiled the possibility of a quantum phase harboring coexisting localized, extended, and critical regions within a one-dimensional quasi-periodic model Wang et al. (2022).
Therefore, we first consider an entirely spin-dependent incommensurate lattice ($M_{z,\uparrow}/M_{z,\downarrow}=-1$) in the Hermitian regime ($\gamma_{\sigma}=0$), and find that a critical phase manifests in this configuration. The phase of this system can be characterized through the mean fractal dimension $\bar{\eta}$, as introduced by Wang et al. (2020):
@@
It is established that, in extended (localized) states, the value of $\bar{\eta}$ tends towards 1 (0), while within the critical phase, $10<\bar{\eta}<$ 1.
In Fig.2(a), the mean fractal dimension is depicted as a function of $M_{z}$ and $t_{so}$, revealing a distinct region where $10<\bar{\eta}<$ 1, indicating the emergence of the critical phase. The phase boundary can be delineated by $M_{z}=2|t_{0}\pm t_{so}|$, where the symbol $+$ signifies the boundary between localization and critical phases, and -- signifies the boundary between extended and critical phases Wang et al. (2020). Fig. 2(b) shows the fractal dimension for various eigenenergies and $M_{z}$ when $t_{so}=0.3$, illustrating that all individual states within the critical region exhibit critical behavior. However, if the incommensurate lattice is entirely spin-independent ($M_{z,\uparrow}/M_{z,\downarrow}=1)$, the critical phase region disappears, as illustrated in Fig. 2(c). Although a region where $\bar{\eta}$ deviates significantly from 1 and 0 can be identified, a further examination of the fractal dimension $\eta$ of the corresponding eigenstates as a function of $M_{z}$ suggests an intermediate regime characterized by the coexistence of extended and localized states rather than a critical phase (Fig.2(d)).
@@
In the pursuit of experimentally detecting localization behavior in an optical lattice system, one common method involves observing the expansion dynamics, as conducted in various studies Billy et al. (2008); Roati et al. (2008); Lüschen et al. (2018). Fig. 3(a) shows simulated time evolutions of wave packets positioned in the extended, critical, and localized phase regions. These simulations are based on an initial wave packet configuration defined as $|{\psi_{j}(t=0)}\rangle=(\sqrt{\pi}a)^{-1/2}e^{-(j-j_{0})^{2}/2a^{2}}|{\uparrow}\rangle$, where $a$ represents the half-width, and the packet is centered at site $j_{0}$. Notably, the outcomes reveal distinctive behaviors: in the extended phase, the atom cloud exhibits ballistic expansion, while in the localized phase, occupancy remains predominantly restricted to the initial site. In the critical phase, a quasi-localization phenomenon is observed, with the atom cloud spreading to neighboring unit cells around the initial site, without fully occupying the entire system.
To further characterize the expansion dynamics, the mean square displacement can be employed as an observable, defined as Wang et al. (2020):

### v2

@@ Abstract
We propose a flexible Raman lattice system for alkaline-earth-like atoms to theoretically investigate localization behaviors in a quasi-periodic lattice with controllable non-Hermiticity. Our analysis demonstrates that critical phases and mobility edges can arise by adjusting spin-dependence of the incommensurate potentials in the Hermitian regime. With non-Hermiticity introduced by spin-selective atom loss, our calculations reveal that critical localization behaviour in this system can be suppressed by dissipation. Our work provides insights into interplay between quasi-periodicity and non-Hermitian physics, offering a new perspective on localization phenomena.
I Introduction
@@
In the presence of a disorder potential, quantum systems can exhibit exponential localization of wave functions induced by wave interference, a phenomenon known as Anderson localization [4]. Ultracold atom systems, with their exceptional purity and high controllability, enable detailed exploration of localization phenomena. Recent experiments have investigated Anderson localization in optical lattices with either random disorder [11] or incommensurate quasi-periodic structures [59, 45, 74].
@@
For ultracold atoms, quasi-periodic optical lattices are created by superimposing two optical lattices with incommensurate wavelengths. In such setups, a notable transition from extended to localized states, commonly referred to as the Anderson transition, can manifest even within one-dimensional configurations [21, 7], contrary to uncorrelated disordered systems where such transitions are expected to occur only in dimensions higher than two according to the scaling theory [66, 1]. These systems also allow for the engineering of mobility edges-boundaries that separate coexisting localized and extended states at discrete energy levels [60, 18, 10, 45, 2, 3, 52]. Recent theoretical studies suggest even more complex behavior, including intermediate critical phases and hybrid phases where localized, extended, and critical states coexist in specific one-dimensional quasi-periodic models [69, 70, 93, 94, 68].
@@
Recent years have seen growing interest in non-Hermitian physics [17, 5] across multiple research domains, including electrical circuits [28, 19], optics and photonics [35, 73, 89, 71], optomechanics [57], NV centers [84, 77], trapped ions [12, 85], superconducting qubits [13] and ultracold atom systems [36, 40, 58]. Non-Hermitian systems display unique features such as exceptional points or rings [88, 58, 42], novel topological invariants, restoration of the dark state [92], and the non-Hermitian skin effect [40, 87, 50]. The Hatano-Nelson model, a fundamental non-Hermitian topological model, combines asymmetric hopping amplitudes with an on-site random potential [26]. However, a comprehensive understanding of the interplay between disorder and non-Hermiticity, particularly the influences of dissipation on localization behavior, remains largely unexplored [79, 80, 25]. Furthermore, experimental methods for studying this relationship between disorder and non-Hermiticity remain limited.
@@
In this paper, we propose a flexible Raman lattice scheme [41] tailored for 173Yb atoms [65, 64, 86] or alkaline-earth metal atoms [39] to theoretically investigate localization behaviors. The Raman lattice configuration induces quasi-periodicity using two optical lattices with incommensurate wavelengths, with the weak lattice potential akin to an incommensurate Zeeman potential. Our model shows that adjusting the frequency of the incommensurate lattice close to the $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition enables spin-dependent control, which can theoretically lead to the emergence of critical phases or mobility edges within the system. We further extend the Raman lattice model with incommensurate potentials into the non-Hermitian regime with dissipation induced by atom loss [87]. Theoretical analysis reveals that dissipation exerts a significant influence on the localization behaviors within the system, providing a platform to model the intricate interplay between disorder potential and non-Hermitian physics.
@@
The subsequent sections of this paper are structured as follows: Sec.II presents the theoretical model Hamiltonian of the flexible Raman lattice and outlines its implementation scheme. In Sec.III-VI, we theoretically analyze the impact of the spin-dependency of the incommensurate lattice and dissipation within the system. Our focus will encompass three distinct scenarios: a completely spin-dependent incommensurate lattice, a partially spin-dependent incommensurate lattice, and an incommensurate lattice coupled with dissipation. Finally, Sec.VII offers a concise summary of the theoretical findings and conclusions derived from this investigation.
II Model Hamiltonian
@@
$H=H_{0}+\sum_{j,\sigma=\uparrow,\downarrow}(\delta_{\sigma}^{j}+i\gamma_{\sigma})n_{j,\sigma}$ (1)
@@
where $\delta^{j}_{\sigma}=M_{z,\sigma}\cos{\left(2\pi\beta j\right)}$, with $M_{z,\sigma}\propto V_{s,\sigma}$ and $\gamma_{\sigma}$ denotes the incommensurate Zeeman potential strength and dissipation strength for spin $\sigma=\uparrow,\downarrow$, respectively. $n_{j,\sigma}=c^{\dagger}_{j,\sigma}c_{j,\sigma}$ is the particle number operator, $\beta$ is an irrational number and:
$H_{0}$ $=-t_{0}\sum_{j,\sigma}\left[\left(c_{j,\uparrow}^{\dagger}c_{j+1,\uparrow}-c_{j,\downarrow}^{\dagger}c_{j+1,\downarrow}\right)+\text{H.c.}\right]$ (2)
$+t_{\text{so}}\sum_{j}\left[\left(c_{j,\uparrow}^{\dagger}c_{j+1,\downarrow}-c_{j,\uparrow}^{\dagger}c_{j-1,\downarrow}\right)+\text{H.c.}\right]$
@@
Here, $t_{0}$ ($t_{so}$) represents the spin-conserved (spin-flip) hopping strength between neighboring sites. The Hamiltonian renders a 1D AIII class topological insulator if replacing the incommensurate Zeeman potential with a constant uniform Zeeman potential [65, 41, 67, 82].
@@
Figure 1: Experimental implementation scheme of critical phase in optical Raman lattice with 173Yb atoms. (a) Experimental setup of the optical Raman lattice consists of a standard 1D optical Raman lattice $E_{1}$, $E_{3}$ with another spin-dependent incommensurate lattice $E_{21}$ and $E_{22}$. (b) Schematic energy diagram with relevant transitions. (c) The incommensurate lattice induces a spin-dependent offset $V_{2,\sigma}(x)$ to the standard optical Raman lattice potential. (d) For the implementation of a perfect spin-dependent incommensurate lattice, the laser detuning should be set to approximately 1.9 GHz blue-detuned from the $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition.
@@
This Hamiltonian can be experimentally implemented through the configuration outlined in Fig. 1(a). To establish the conventional optical Raman lattice potential [41, 72, 65, 83, 33], a standing-wave beam $E_{1}$ with $y$ polarization and a free-running beam $E_{3}$ with z polarization are employed. The single-photon detuning of these two beams can be finely tuned to approximately 11 GHz blue-detuned from the $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, as depicted in Fig. 1(b). In this setup, the primary lattice, characterized by a wave vector of $k_{0}=2\pi/\lambda$, exhibits nearly spin-independent lattice potential $V_{1,\sigma}(x)=V_{p,\sigma}\cos^{2}{k_{0}x}$ with $V_{p,\uparrow}/V_{p,\downarrow}\sim 1.01$, where $V_{p,\uparrow}/V_{p,\downarrow}$ denotes the primary lattice depth for spin up or spin down atoms. In the subsequent discussion, we will neglect the spin-dependent effects of the primary lattice potential ($V_{p,\uparrow}=V_{p,\downarrow}=V_{p}$). Moreover, two additional beams propagating at an angle of $\pm$52 degrees relative to the x direction are utilized to induce a spin-dependent incommensurate lattice featuring a wave vector of $k_{1}=k_{0}\cos{\theta}$, giving rise to a secondary incommensurate Zeeman potential $V_{2,\sigma}=V_{s,\sigma}\cos^{2}{k_{1}x}$ with $V_{s,\sigma}$ representing the incommensurate lattice depth, as illustrated in Fig. 1(c). The spin-dependence of this incommensurate lattice can be controlled through the detuning $\Delta$ of this lattice beam with respect to the $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition. When $\Delta\sim 1.9~\text{GHz}$, the incommensurate lattice becomes entirely spin-dependent with $M_{z,\uparrow}/M_{z,\downarrow}=-1$ (Fig. 1(d)). To achieve a complete spin-independent configuration, the detuning of the incommensurate lattice can be set larger than 13GHz. The Raman coupling in this setup can be denoted by $\mathcal{M}(x)=M_{R}\cos{k_{0}x}$. The introduction of dissipation into this Raman lattice setup can be realized through adding a nearly resonant loss beam with wavelength close to $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition [87].
@@
Here, $\frac{\hbar^{2}k_{x}^{2}}{2m}$ is the kinetic energy term and $\delta$ denotes the two-photon detuning. The incommensurate lattice introduces a weak spin-dependent energy offset characterized by the irrational number $\beta=\cos\theta\approx 0.6157$, where $\beta$ is determined by the tilted angle. In our proposed configuration, which employs $m_{F}$ states within the same hyperfine manifold as $|{\uparrow,\downarrow}\rangle$, the spin-dependence of the incommensurate lattice can be independently tuned by manipulating the detuning of the lattice beam. This independent controllability is advantageous, as it becomes challenging to simultaneously tune $\beta$ and maintain a spin-dependent incommensurate lattice when $\beta$ is set by the laser wavelength ratio. As a result, our approach significantly broadens the range of accessible $\beta$ values.
Notably, our method differs from existing schemes reported in [69, 70, 94]. Those prior proposals were designed for alkali-metal atoms and rely on counter-propagating laser beams with wavelengths near the D1 and D2 transitions, rendering such schemes incompatible with alkaline-earth atoms due to the lack of these transitions [27]. In those setups, the incommensurability parameter $\beta$ is determined by the ratio of the laser wavelengths, inherently restricting it to a narrow, specific range. Furthermore, spin dependence of the incommensurate lattice is controlled via the polarization of the laser light, which requires the spin states to be encoded in different hyperfine states. This contrasts with our configuration and previously demonstrated experimental realizations, where spin states are defined using different $m_{F}$ states within the same hyperfine manifold. Our scheme is particularly suited for fermionic alkaline-earth atoms such as 173Yb, a system already demonstrated in optical Raman lattices [65, 64, 87]. The fermionic nature of 173Yb, together with the potential SU(N) symmetry of alkaline-earth atoms, can lead to phenomena distinct from bosonic systems, especially in the many-body regime [51, 68, 69].
III Localization behavior in the Hermitian regime
Except for the localized phase, recent theoretical investigations have predicted the existence of a third fundamental phase, called critical phases, positioned between the localized and extended phases within such a spin-orbit coupled optical lattice featuring an incommensurate Zeeman potential [69, 65, 87, 94]. Noteworthy features of these critical phases encompass critical spectral statistics [20, 46, 9], multifractal properties of wave functions [23, 48, 15], and dynamical evolutions [29, 31, 34]. Moreover, recent theoretical investigations have also unveiled the possibility of a quantum phase harboring coexisting localized, extended, and critical regions within a one-dimensional quasi-periodic model [70].
Therefore, we first consider an entirely spin-dependent incommensurate lattice ($M_{z,\uparrow}/M_{z,\downarrow}=-1$) in the Hermitian regime ($\gamma_{\sigma}=0$), and find that a critical phase manifests in this configuration. The phase of this system can be characterized through the mean fractal dimension $\bar{\eta}$, as introduced by [69]:
@@ III Localization behavior in the Hermitian regime
It is established that, in extended (localized) states, the value of $\bar{\eta}$ tends towards 1 (0), while within the critical phase, $0<\bar{\eta}<1$.
In Fig.2(a), the mean fractal dimension is depicted as a function of $M_{z}$ and $t_{so}$, revealing a distinct region where $0<\bar{\eta}<1$, indicating the emergence of the critical phase. The phase boundary can be delineated by $M_{z}=2|t_{0}\pm t_{so}|$, where the symbol ++ signifies the boundary between localization and critical phases, and - signifies the boundary between extended and critical phases [69]. Fig. 2(b) shows the fractal dimension for various eigenenergies and $M_{z}$ when $t_{so}=0.3$, illustrating that all individual states within the critical region exhibit critical behavior. However, if the incommensurate lattice is entirely spin-independent ($M_{z,\uparrow}/M_{z,\downarrow}=1)$, the critical phase region disappears, as illustrated in Fig. 2(c). Although a region where $\bar{\eta}$ deviates significantly from 1 and 0 can be identified, a further examination of the fractal dimension $\eta$ of the corresponding eigenstates as a function of $M_{z}$ suggests an intermediate regime characterized by the coexistence of extended and localized states rather than a critical phase (Fig.2(d)).
@@
In the pursuit of experimentally detecting localization behavior in an optical lattice system, one common method involves observing the expansion dynamics, as conducted in various studies [11, 59, 45]. Fig. 3(a) shows simulated time evolutions (in units of $\hbar/t_{0}$) of wave packets positioned in the extended, critical, and localized phase regions. These simulations are based on an initial wave packet configuration defined as $|{\psi_{j}(t=0)}\rangle=(\sqrt{\pi}a)^{-1/2}e^{-(j-j_{0})^{2}/2a^{2}}|{\uparrow}\rangle$, where $a$ represents the half-width, and the packet is centered at site $j_{0}$. Notably, the outcomes reveal distinctive behaviors: in the extended phase, the atom cloud exhibits ballistic expansion, while in the localized phase, occupancy remains predominantly restricted to the initial site. In the critical phase, a quasi-localization phenomenon is observed, with the atom cloud spreading to neighboring unit cells around the initial site, without fully occupying the entire system.
To further characterize the expansion dynamics, the mean square displacement can be employed as an observable, defined as [69]: